# Vision-driven River Following of UAV via Safe Reinforcement Learning using Semantic Dynamics Model

### v1

@@ Abstract
Vision-driven autonomous river following by Unmanned Aerial Vehicles is critical for applications such as rescue, surveillance, and environmental monitoring, particularly in dense riverine environments where GPS signals are unreliable. We formalize river following as a coverage control problem in which the reward function is submodular—yielding diminishing returns as more unique river segments are visited—thereby framing the task as a Submodular Markov Decision Process. Building on the Safe Reinforcement Learning (SafeRL) paradigm and the First Order Constrained Optimization in Policy Space algorithm, we propose three key contributions. First, we introduce Marginal Gain Advantage Estimation (MGAE), which refines the reward advantage function by using a sliding-window baseline computed from historical episodic returns, thus aligning the advantage estimation with the agent’s evolving recognition of action value in non-Markovian settings. Second, we develop a Semantic Dynamics Model (SDM) based on patchified water semantic masks that provides more interpretable and data-efficient short-term prediction of future observations compared to latent vision dynamics models. Third, we present the Constrained Actor Dynamics Estimator (CADE) architecture, which integrates the actor, the cost estimator, and SDM for cost advantage estimation to form a model-based SafeRL framework capable of solving partially observable Constrained Submodular Markov Decision Processes. Simulation results demonstrate that MGAE achieves faster convergence and superior performance over traditional critic-based methods like Generalized Advantage Estimation. SDM provides more accurate short-term state predictions that enable the cost estimator to better predict potential violations. Overall, CADE effectively integrates safety regulation into model-based RL, with the Lagrangian approach achieving the "soft" balance of reward and safety during training, while the safety layer enhances performance during inference by "hard" action overlay. Our code is available on Github (https://github.com/EdisonPricehan/omnisafe-cade/tree/cacd).
@@
Compare Lagrangian training with cost-planning safety layer for safe policy execution.
@@ 1 INTRODUCTION
Autonomous navigation along riverine environments is a crucial task for Unmanned Aerial Vehicles (UAVs) in applications such as environmental monitoring, surveillance and rescue missions. While previous work has focused on vision-driven navigation using segmentation masks for water detection [1], these approaches often rely on manual tuning and simplistic environmental models, making them unsuitable for complex, real-world scenarios involving obstacles like trees, bridges, and varying river geometries [2, 3]. Moreover, conventional GPS-based waypoint navigation suffers from inaccuracies due to seasonal variations in river shape and can be further compromised in dense jungle riverine environments where GPS signals are unreliable or unavailable, making these methods unsuitable for safe and precise river following missions. Thus, vision-based navigation becomes essential in such environments [4].
@@
Vision-driven river following, as a specific form of coverage control problem, poses significant challenges for Reinforcement Learning (RL) due to its partial observability, submodular reward structure, and non-Markovian nature. The partial observability arises from the agent’s reliance on vision-based observations, which provide only a limited, first-person view of the environment, making it difficult to perceive the entire river layout [4]. Unlike standard Markov Decision Processes (MDPs, [5]), where the reward depends solely on the current state and action, river-following rewards are non-Markovian, as they depend on the agent’s historical trajectory and are given only when new, unvisited segments of the river are explored [4]. Furthermore, this reward function is submodular, meaning that the marginal gain from visiting river segments diminishes as more segments are explored, akin to a diminishing returns property in optimization. This submodular structure, combined with the fact that rewards depend on the agent’s past trajectory rather than just the current state, renders the task a Submodular Markov Decision Process (SMDP, [6]). As a result, a more sophisticated advantage estimation method is required to effectively leverage the submodular nature of the reward and the dependence on the trajectory. Furthermore, safety is a critical concern in such environments, necessitating the use of Constrained Markov Decision Processes (CMDPs, [7]) that introduce a cost function $C$ for penalizing unsafe actions, extending MDPs with the necessary safety constraints. These challenges necessitate more sophisticated reward estimation and policy optimization methods that can effectively capture temporal dependencies, leverage the submodular reward structure, and address safety requirements, leading to the formulation of a Constrained Submodular Markov Decision Process (CSMDP).
@@
To address the submodular and non-Markovian nature of the reward function in river following tasks, we propose a Marginal Gain Advantage Estimation (MGAE) method. Unlike traditional advantage estimators such as Generalized Advantage Estimation (GAE, [8]), which look forward along trajectory and rely on a state-dependent value function as the baseline, MGAE combines both forward and backward perspectives. It looks forward to accumulate actual rewards while looking backward into the agent’s trajectory to estimate cumulative marginal gains. This approach captures the historical dependencies essential in environments where rewards are given only for visiting new, unvisited river segments, resulting in a less biased estimation of the agent’s true progress. By accounting for the submodular reward structure and using the average episodic return as a baseline, MGAE is particularly well-suited for tasks where the reward depends on the agent’s exploration history, rather than solely on the current state or observation.
@@
In addition to improving the reward advantage estimation in SMDP, we introduce a Semantic Dynamics Model (SDM) to predict future states based on the agent’s visual input. In the context of visual semantic navigation [9, 10, 11, 12], some model-based safe reinforcement learning approaches leverage semantic information to guide decision-making without explicit knowledge of spatial context. SDM complements this by exploiting the dynamics of visual observations in both geometric and semantic manners, enhancing the agent’s understanding of the environment. SDM explicitly incorporates homography transformations between observations in first-person view, preserving both geometric and semantic information. Unlike state-of-the-art latent dynamics models [13, 14, 15, 16, 17], which abstract away raw pixel data into latent vectors—losing essential spatial and semantic relationships—our approach offers greater interpretability by modeling state transitions based on the actual geometric structure of the riverine environment. This makes our model particularly suitable for real-world applications like outdoor navigation, where spatial reasoning and explainability are critical. SDM helps the agent anticipate potential hazards to safely navigate through complex riverine environments. By integrating this model with our advantage estimation method, the agent can make more informed decisions that account for both future rewards and potential safety violations, leading to safer exploration and better task performance.
@@
Our architecture, Constrained Actor Dynamics Estimator (CADE), brings together MGAE and SDM within the framework of safe reinforcement learning. The CADE architecture is designed to optimize a policy that balances reward maximization with safety constraints, ensuring that the UAV navigates effectively while avoiding collisions or other hazardous behaviors. Inspired by recent advances in constrained policy optimization methods [18, 19], our approach applies the Lagrangian relaxation technique to handle safety constraints during training, making it well-suited for real-world deployment in safety-critical environments. This work builds upon and extends our prior research in vision-based navigation [4] using safe reinforcement learning [20], addressing the unique challenges of autonomous river following in complex environments with only visual inputs.
@@
    We develop the Semantic Dynamics Model (SDM) based on homography transformation between adjacent semantic observations, to predict future observations and help improve safer decision-making for vision-driven navigation.
@@
This paper is organized as follows: Section 2 introduces related work in terms of river following and vision-driven model-based safe reinforcement learning that solves this coverage control problem. Section 3 explains our contributions in marginal gain advantage estimation for reward, semantic dynamics model, cost advantage from the cost estimator, and the combined constrained actor dynamics estimator architecture. Section 4 shows our experiment design, and Section 5 shows the comparison results and analyses. Section 6 discusses several design choices and their implications for future work. Finally, Section 7 concludes our paper with key contributions and conclusions.
@@ 2 RELATED WORK
This section introduces the related work from the viewpoints of several perspectives for vision-driven safe river following task: existing approaches and environments for autonomous semantic perception driven river following (Section 2.1), reward advantage estimation methods in usual Markov Decision Processes (Section 2.2), vision representation learning and vision-based dynamics model construction (Section 2.3), and methods using dynamics model in safe reinforcement learning (Section 2.4).
@@ 2.1 River Navigation
Autonomous UAV navigation along curved river segments has previously been implemented using HSV (hue, saturation, and value) threshold-based water segmentation and a proportional heading controller [1]. However, the semantic mask extraction in this approach requires precise manual tuning, which may limit its adaptability to varying river types and lighting conditions. Furthermore, the proportional controller struggles with navigation in rivers with tributaries, while a fixed-altitude flight path can pose challenges when encountering bridges. Recent advances in image segmentation, particularly deep convolutional neural networks, have enabled robust and generalizable semantic segmentation for robotics and UAV navigation tasks, far surpassing classic HSV thresholding methods that require extensive manual tuning and are sensitive to environmental variations. Modern approaches leverage semantic segmentation masks as input to control policies, facilitating robust navigation and obstacle avoidance in complex environments [21, 22, 23, 24, 25, 26]. For example, [26] demonstrates that a segmentation-based preprocessing step (GA-Nav) classifies terrain navigability in unstructured outdoor environments, leading to a $10\%$ increase in navigation success and a $4.6\%$ to $13.9\%$ decrease in trajectory roughness compared to RGB-only policies on a ground vehicle. Overall, semantic masks serve as domain-invariant, high-level inputs that boost sim-to-real transferability and reduce sample complexity compared to raw RGB or autoencoder embeddings. Our approach builds on this trend by employing patchified semantic water masks as the observation space (Section 3.3), which enhances sample efficiency and robustness to visual noise.
Researchers employed imitation learning to develop a river navigation policy based on vision input in an Unreal Engine-based riverine environment [27, 28]. They shown that separate training of the image encoder (variational autoencoder in their experiments) and the control policy is superior to an end-to-end neural network controller in terms of navigation distance and intervention rate. However, their method has several limitations that impact its applicability. The intervention-based DAgger [29] algorithm depends on manual guidance and reset criteria during training to ensure safety, which is human involved for safety feedback. Additionally, it restricts the agent’s actions to a 2D plane, lacking the full 3D maneuverability required for UAV tasks. The absence of public access to the environment further limits reproducibility and broader research collaboration. These constraints make the method heavily dependent on human involvement during training, for both task completion and obstacle avoidance.
@@ 2.2 Advantage Estimation
One of the most widely adopted methods is Generalized Advantage Estimation (GAE, [8]) in the Proximal Policy Optimization (PPO) framework. GAE calculates the advantage function by blending TD errors across multiple time steps with a smoothing parameter $\lambda$, which provides a trade-off between bias and variance. By leveraging this temporal smoothing, GAE is particularly effective in scenarios where accurate, low-variance advantage estimates are essential for stable policy learning. Since GAE downgrades to TD when $\lambda=0$, we set $\lambda=0.95$ for all GAE-related experiments.
For partially observable or non-Markovian reward environments, V-trace is an advantage estimation method used within the IMPALA architecture [33]. V-trace corrects off-policy trajectories by scaling the advantage estimates according to importance sampling weights, reducing variance introduced by sampling errors and ensuring stable advantage estimation in large-scale distributed setups.
While these methods have demonstrated success across various domains, they rely on state-wise advantage estimation, focusing primarily on short-term TD signals. However, this can be limiting in environments with non-Markovian and submodular reward structures, where the immediate reward signal is insufficient for accurately reflecting long-term progress. Such settings, including river or track-following tasks, require the agent to maintain awareness of visited regions to avoid revisiting and to maximize task-specific rewards. Traditional state-wise approaches often fail to capture this trajectory-level context, as they overlook the cumulative effect of prior actions on the current reward.
To address these limitations, we propose Marginal Gain Advantage Estimation (MGAE), which incorporates the cumulative marginal gain over trajectories to reflect the agent’s progress in visiting novel states. Unlike GAE and other state-wise methods, MGAE tracks the historical dependencies of rewards by estimating advantage based on the accumulated marginal gains, making it better suited for non-Markovian tasks with history-dependent rewards. This approach provides a less biased and more informative signal for policy optimization in complex navigation tasks, where maintaining trajectory-level awareness is critical for effective and efficient exploration.
@@ 2.3 Vision Dynamics Model
In recent years, model-based reinforcement learning approaches, such as PlaNet [15] and the Dreamer series [13, 16, 14], have demonstrated the potential of learning latent dynamics models directly from high-dimensional visual inputs. PlaNet introduced a recurrent state-space model (RSSM) for planning in latent space. DreamerV1 [13] extended this by using a variational autoencoder (VAE) [34] to encode raw RGB observations into a compact latent space, enabling efficient simulation and decision-making. DreamerV2 [16] improved upon DreamerV1 with better stability and scalability, supporting longer training horizons and continuous control tasks. DreamerV3 [14] further refined the framework by enhancing generalization and planning efficiency in complex environments, optimizing both reward prediction and policy learning within the latent dynamics model.
While PlaNet and Dreamer models have advanced the learning of compact latent dynamics, they lack interpretability, overlook explicit geometrical information, and do not emphasize task-relevant semantics, potentially leading to inefficiencies in spatially complex environments. Additionally, if an agent needs to learn a non-Markovian cost function based on vision input, a latent vector input may obscure dependencies and make learning less intuitive. In contrast, the proposed semantic dynamics model (SDM) integrates interpretable, geometrically structured representations and semantically meaningful features, enhancing the agent’s capacity for efficient learning, spatial coherence, and decision-making in complex, task-oriented environments.
Deep Visual Foresight [35] provides a flexible framework for robot motion planning through self-supervised, pixel-based prediction, enabling online adaptation without semantic input. While highly adaptable, this approach can face challenges with substantial lighting changes or unfamiliar terrain and incurs significant computational cost due to raw pixel processing. The SDM, by contrast, reduces visual dimensionality through a patchified semantic mask, yielding a compact, interpretable representation with essential environmental context preserved. Though SDM uses 2D homography rather than full 3D geometry, it captures spatial transformations efficiently, offering a structured and computationally efficient model for geometrically constrained ego-centric vision-driven tasks.

### v2

@@ Abstract
Vision-driven autonomous river following by Unmanned Aerial Vehicles is critical for applications such as rescue, surveillance, and environmental monitoring, particularly in dense riverine environments where GPS signals are unreliable. These safety-critical navigation tasks must satisfy hard safety constraints while optimizing performance. Moreover, the reward in river following is inherently history-dependent (non-Markovian) by which river segment has already been visited, making it challenging for standard safe Reinforcement Learning (SafeRL). To address these gaps, we cast river following as a coverage control problem with a submodular reward that exhibits diminishing returns as more river segments are visited, framing the task as a Submodular Markov Decision Process. Building on the SafeRL paradigm and the First Order Constrained Optimization in Policy Space (FOCOPS) algorithm, we propose three contributions.
First, we introduce the Marginal Gain Advantage Estimation (MGAE), which refines the reward advantage function using a sliding-window baseline calculated from historical episodic returns, aligning the advantage estimate with non-Markovian dynamics. Second, we develop a Semantic Dynamics Model (SDM) based on patchified water semantic masks, offering more interpretable and data-efficient short-term prediction of future observations compared to latent vision dynamics models. Third, we present the Constrained Actor Dynamics Estimator (CADE) architecture, which integrates the actor, cost estimator, and SDM for cost advantage estimation to form a model-based SafeRL framework capable of solving partially observable Constrained Submodular Markov Decision Processes.
The simulation results demonstrate that MGAE achieves faster convergence and superior performance compared to critic-based methods like Generalized Advantage Estimation. SDM provides more accurate short-term state predictions, enabling the cost estimator to better predict potential violations. Overall, CADE effectively integrates safety regulation into model-based RL, with the Lagrangian approach providing a "soft" balance between reward and safety during training, while the safety layer enhances inference by imposing a "hard" action overlay. Our code is publicly available on Github (https://github.com/EdisonPricehan/omnisafe-cade/tree/cacd).
@@
Compare Lagrangian training with the cost-planning safety layer for safe policy execution.
@@ 1 INTRODUCTION
Autonomous navigation in riverine environments is a crucial task for Unmanned Aerial Vehicles (UAVs) in applications such as environmental monitoring, surveillance, and rescue missions. Although previous work has focused on vision-driven navigation using segmentation masks for water detection [1], these approaches often rely on manual tuning and simplistic environmental models. Manually tuned controller can work in controlled settings, but as environmental variability (trees, bridges, and river geometries) increases it becomes brittle to distribution shift and costly, subjective, and time-consuming to retune for each site [2, 3]. In contrast, a common alternative is GPS-based waypoint navigation, which tracks preplanned river centerlines based on prior maps. However, seasonal changes in channel shape and canopy-induced GPS outages degrade these plans, undermining the reliability of safe and precise river following. Thus, vision-based navigation becomes essential in such environments [4].
@@
Vision-driven river following, as a specific form of coverage control problem, poses significant challenges to Reinforcement Learning (RL) due to its partial observability, submodular reward structure, and non-Markovian nature. The partial observability arises from the agent’s reliance on vision-based observations, which provide only a limited first-person view of the environment, making it difficult to perceive the entire river layout [4]. Unlike standard Markov Decision Processes (MDPs, [5]), where the reward depends solely on the current state and action, river-following rewards are non-Markovian, as they depend on the agent’s historical trajectory and are given only when new unvisited segments of the river are explored [4]. History-dependent rewards also arise in information gathering problems, where rewards encode reductions in uncertainty, making the return depend on the observation history rather than only the current state [6]. Furthermore, this reward function is submodular, meaning that the marginal gain from visiting river segments diminishes as more segments are explored, akin to a diminishing returns property in optimization. This submodular structure, combined with the fact that the rewards depend on the agent’s past trajectory rather than just the current state, makes the task a Submodular Markov Decision Process (SMDP, [7]). As a result, a more sophisticated advantage estimation method is required to effectively leverage the submodular nature of the reward and the dependence on the trajectory. Furthermore, safety is a critical concern in such environments, necessitating the use of Constrained Markov Decision Processes (CMDPs, [8]) that introduce a cost function $C$ to penalize unsafe actions, extending MDPs with the necessary safety constraints. These challenges require more sophisticated reward estimation and policy optimization methods that can effectively capture temporal dependencies, leverage the submodular reward structure, and address safety requirements, leading to the formulation of a Constrained Submodular Markov Decision Process (CSMDP).
@@
To address the submodular and non-Markovian nature of the reward function in river following tasks, we propose a Marginal Gain Advantage Estimation (MGAE) method. Unlike traditional advantage estimators such as Generalized Advantage Estimation (GAE, [9]), which look forward along the trajectory and rely on a state-dependent value function as the baseline, MGAE combines both forward and backward perspectives. It looks forward to accumulate actual rewards while looking backward into the agent’s trajectory to estimate cumulative marginal gains. This approach captures the historical dependencies essential in environments where rewards are given only for visiting new, unvisited river segments, resulting in a less biased estimation of the agent’s true progress. By accounting for the submodular reward structure and using the average episodic return as a baseline, MGAE is particularly well-suited for tasks where the reward depends on the agent’s exploration history, rather than solely on the current state or observation.
@@
In addition to improving the reward advantage estimation in SMDP, we introduce a Semantic Dynamics Model (SDM) to predict future states based on the agent’s visual input. In the context of visual semantic navigation [10, 11, 12, 13], some model-based safe reinforcement learning approaches leverage semantic information to guide decision-making without explicit knowledge of spatial context. SDM complements this by exploiting the dynamics of visual observations in both geometric and semantic manners, enhancing the agent’s understanding of the environment. SDM explicitly incorporates homography transformations between observations in first-person view, preserving both geometric and semantic information. Unlike state-of-the-art latent dynamics models [14, 15, 16, 17, 18], which abstract raw pixel data into latent vectors, losing essential spatial and semantic relationships, our approach offers greater interpretability by modeling state transitions based on the actual geometric structure of the riverine environment. This makes our model particularly suitable for real-world applications, such as outdoor navigation, where spatial reasoning and explainability are critical. SDM helps the agent anticipate potential hazards to safely navigate through complex riverine environments. By integrating this model with our advantage estimation method, the agent can make more informed decisions that account for both future rewards and potential safety violations, leading to safer exploration and better task performance.
@@
Our architecture, Constrained Actor Dynamics Estimator (CADE), brings together MGAE and SDM within the framework of safe reinforcement learning. The CADE architecture is designed to optimize a policy that balances reward maximization with safety constraints, ensuring that the UAV navigates effectively while avoiding collisions or other hazardous behaviors. Inspired by recent advances in constrained policy optimization methods [19, 20], our approach applies the Lagrangian relaxation technique to handle safety constraints during training, making it suitable for real-world deployment in safety-critical environments. This work builds upon and extends our prior research in vision-based navigation [4] using safe reinforcement learning [21], addressing the unique challenges of autonomous river following in complex environments with only visual inputs.
@@
    We develop the Semantic Dynamics Model (SDM) based on homography transformation between adjacent semantic observations to predict future observations and help improve safer decision-making for vision-driven navigation.
@@
This paper is organized as follows: Section 2 introduces related work in terms of river following and vision-driven model-based safe reinforcement learning that solves this coverage control problem. Section 3 explains our contributions in marginal gain advantage estimation for reward, the semantic dynamics model, the cost advantage of the cost estimator, and the combined architecture of the constrained actor dynamics estimator architecture. Section 4 shows our experiment design, and Section 5 shows the comparison results and analyses. Section 6 discusses several design choices and their implications for future work. Finally, Section 7 concludes our paper with key contributions and conclusions.
@@ 2 RELATED WORK
This section introduces the related work from the viewpoints of several perspectives for the vision-driven safe river following task: existing approaches and environments for autonomous semantic perception driven river following (Section 2.1), reward advantage estimation methods in usual Markov Decision Processes (Section 2.2), vision representation learning and vision-based dynamics model construction (Section 2.3), and methods using the dynamics model in safe reinforcement learning (Section 2.4).
@@ 2.1 River Navigation
Autonomous UAV navigation along curved river segments has previously been implemented using threshold-based HSV (hue, saturation, and value) water segmentation and a proportional heading controller [1]. However, the semantic mask extraction in this approach requires a precise manual tuning, which can limit its adaptability to varying river types and lighting conditions. Furthermore, the proportional controller struggles with navigation in rivers with tributaries, while a fixed-altitude flight path can pose challenges when encountering bridges. Recent advances in image segmentation, particularly deep convolutional neural networks, have enabled robust and generalizable semantic segmentation for robotics and UAV navigation tasks, far surpassing classic HSV thresholding methods that require extensive manual tuning and are sensitive to environmental variations. Modern approaches leverage semantic segmentation masks as input to control policies, facilitating robust navigation and obstacle avoidance in complex environments [22, 23, 24, 25, 26, 27]. For example, [27] shows that a segmentation-based preprocessing step (GA-Nav) classifies terrain navigability in unstructured outdoor environments, leading to a $10\%$ increase in navigation success and a $4.6\%$ to $13.9\%$ decrease in trajectory roughness compared to RGB-only policies on a ground vehicle. In general, semantic masks serve as domain-invariant, high-level inputs that boost sim-to-real transferability and reduce sample complexity compared to raw RGB or autoencoder embeddings. Our approach builds on this trend by employing patchified semantic water masks as the observation space (Section 3.3), which enhances sample efficiency and the robustness to visual noise.
The researchers used imitation learning to develop a river navigation policy based on vision input in an environment based on Unreal Engine [28, 29]. They showed that separate training of the image encoder (variational autoencoder in their experiments) and the control policy is superior to an end-to-end neural network controller in terms of navigation distance and intervention rate. However, their method has several limitations that impact its applicability. The intervention-based DAgger algorithm [30] depends on manual guidance and reset criteria during training to ensure safety, which is human involved for safety feedback. In addition, it restricts the agent’s actions to a 2D plane, lacking the full 3D maneuverability required for UAV tasks. The lack of public access to the environment further limits reproducibility and broader research collaboration. These constraints make the method highly dependent on human involvement during training, for both task completion and obstacle avoidance.
@@ 2.2 Advantage Estimation
One of the most widely adopted methods is Generalized Advantage Estimation (GAE, [9]) in the Proximal Policy Optimization (PPO) framework. GAE calculates the advantage function by blending TD errors in multiple time steps with a smoothing parameter $\lambda$, which provides a trade-off between bias and variance. Using this temporal smoothing, GAE is particularly effective in scenarios where accurate, low-variance advantage estimates are essential for stable policy learning. Since GAE downgrades to TD when $\lambda=0$, we set $\lambda=0.95$ for all GAE-related experiments.
For partially observable or non-Markovian reward environments, V-trace is an advantage estimation method used within the IMPALA architecture [34]. V-trace corrects off-policy trajectories by scaling the advantage estimates according to importance sampling weights, reducing variance introduced by sampling errors, and ensuring stable advantage estimation in large-scale distributed setups.
Although these methods have shown success across various domains, they rely on state-wise advantage estimation, focusing primarily on short-term TD signals. However, this can be limiting in environments with non-Markovian and submodular reward structures, where the immediate reward signal is insufficient to accurately reflect long-term progress. Such settings, including river or track-following tasks, require the agent to maintain awareness of visited regions to avoid revisiting and to maximize task-specific rewards. Traditional state-wise approaches often fail to capture this trajectory-level context, as they overlook the cumulative effect of prior actions on the current reward.
To address these limitations, we propose the Marginal Gain Advantage Estimation (MGAE), which incorporates the cumulative marginal gain over trajectories to reflect the agent’s progress in visiting new states. Unlike GAE and other state-wise methods, MGAE tracks the historical dependencies of rewards by estimating advantage based on the accumulated marginal gains, making it better suited for non-Markovian tasks with history-dependent rewards. This approach provides a less biased and more informative signal for policy optimization in complex navigation tasks, where maintaining trajectory-level awareness is critical for effective and efficient exploration.
@@ 2.3 Vision Dynamics Model
In recent years, model-based reinforcement learning approaches, such as PlaNet [16] and the Dreamer series [14, 17, 15], have demonstrated the potential of learning latent dynamics models directly from high-dimensional visual inputs. PlaNet introduced a recurrent state-space model (RSSM) for planning in latent space. DreamerV1 [14] extended this by using a variational autoencoder (VAE) [35] to encode raw RGB observations into a compact latent space, enabling efficient simulation and decision-making. DreamerV2 [17] improved DreamerV1 with better stability and scalability, supporting longer training horizons and continuous control tasks. DreamerV3 [15] further refined the framework by improving generalization and planning efficiency in complex environments, optimizing both reward prediction and policy learning within the latent dynamics model.
Different levels of abstraction in vision dynamics models depend on the task and environment [36]. While PlaNet and Dreamer models have advanced the learning of compact latent dynamics, they lack interpretability, overlook explicit geometrical information, and do not emphasize task-relevant semantics, potentially leading to efficiency, object permanence and temporal consistency issues in riverine environments. Additionally, if an agent needs to learn a Markovian cost function based on vision input, a latent vector input may obscure dependencies and make learning less intuitive. In contrast, the proposed semantic dynamics model (SDM) integrates interpretable geometrically structured representations and semantically meaningful features, enhancing the agent’s capacity for efficient learning, spatial coherence and decision-making in complex task-oriented environments.
Deep Visual Foresight [37] provides a flexible framework for robot motion planning through pixel-based self-supervised prediction, enabling online adaptation without semantic input. While highly adaptable, this approach can face challenges with substantial lighting changes or unfamiliar terrain and incurs significant computational cost due to raw pixel processing. The SDM, in contrast, reduces visual dimensionality through a patchified semantic mask, yielding a compact, interpretable representation with essential environmental context preserved. Although SDM uses 2D homography rather than full 3D geometry, it captures spatial transformations efficiently, offering a structured and computationally efficient model for geometrically constrained ego-centric vision-driven tasks.